# Tiny but Mighty: A Software-Hardware Co-Design Approach for Efficient Multimodal Inference on Battery-Powered Small Devices

### v2

@@ Abstract
Large Multimodal Models (LMMs) are inherently modular, consisting of vision and audio encoders, projectors, and large language models. Yet, they are almost always executed monolithically, which underutilizes the heterogeneous accelerators (NPUs, GPUs, DSPs) in modern SoCs and leads to high end-to-end latency. In this paper, we present Nanomind, a hardware–software co-design inference framework for Large Multimodal Models (LMMs) that breaks large models into modular “bricks” (vision, language, audio, etc.) and maps each to its ideal accelerator. The key insight is that large models can be broken into modular components and scheduled to run on the most appropriate compute units. It performs module-level dynamic offloading across accelerators on unified-memory SoCs. By combining customized hardware design, system-level scheduling, and optimized low-bit computation kernels, we demonstrate our framework with a compact, battery-powered device capable of running LMMs entirely on-device. This prototype functions as a self-contained intelligent assistant that requires no network connectivity, while achieving higher throughput and superior power efficiency under strict resource constraints. The design further bypasses CPU bottlenecks and reduces redundant memory usage through token-aware buffer management and module-level coordination. Our system outperforms existing implementations in resource efficiency, cutting energy consumption by 42.3% and GPU memory usage by 11.2%. This enables a battery-powered device to run LlaVA-OneVision with a camera for nearly half a day and LLaMA-3-8B for voice interactions up to almost 20.8 hours.
@@ 1 Introduction
To overcome these challenges, we introduce Nanomind, the first fully on-device inference framework that partitions large multimodal models into modular, independently executable components and dynamically assigns each to its optimal compute unit—GPU, NPU, or CPU. Built on a tightly integrated software–hardware co-design, we demonstrated Nanomind by developing a small battery-powered device, as shown in Figure 10. With this custom-designed hardware, our system outperforms mainstream frameworks running on existing commodity off-the-shelf platforms. We also design an event-driven On-Demand Cascade Inference Pipeline 2. Only the minimal output needed—such as a text string or an embedding vector—is retained and passed to the next stage. This results in a lightweight, domino-like chain of execution.
@@
  • Custom Hardware–Software Co-Design. On the hardware side, we use a commodity RK3566 SoC with integrated GPU and NPU, maximized memory bandwidth with four parallel LPDDR4x modules, and add a dedicated power management unit (PMU) for real-time energy monitoring. On the software side, we implement custom 2-bit, 4-bit, and 8-bit GEMM kernels tailored to our hardware, along with an offloading scheduler and drivers to accelerate quantized tensor operations on both GPU and NPU.
@@
By employing these efforts, a tiny device can efficiently operate LLMs and LMMs (LlaVa Liu et al. (2023b; a), Qwen-VL series Bai et al. (2023); Wang et al. (2024b)) within constrained hardware resources by directly offloading workloads to the on-device GPU or NPU based on power and memory usage, bypassing CPU operation time. This approach enhances inference performance and significantly reduces power consumption. Our work lays the foundation for bringing LLMs to resource-constrained environments, enabling the development of responsive, power-efficient, and intelligent systems. It opens the door to democratizing LMM deployment on small devices, transforming how we interact with AI in everyday settings.
@@ 2.2 On-Device Inference Systems and Frameworks
Inefficiencies in llama.cpp While llama.cpp introduces a layer-wise offloading mechanism, its workload distribution is inefficient for small devices, especially on modern unified memory platforms. Although it allows computation to be split between the CPU and GPU, GPU execution still relies on CPU-controlled data transfers via buffers, increasing memory overhead during inference. Figure 9 in the Appendix illustrates this offloading workflow. When a GPU is available, tensors can be assigned the GGML_BACKEND_GPU flag, allowing ggml_compute_forward() to offload computations to the GPU. This involves transferring key tensors from CPU memory, while the CPU must continuously write to buffers and maintain separate memory allocation, leading to additional overhead. This type of framework enables LLM deployment on edge devices but follows server-side designs with separate CPU and GPU memory. In contrast, modern edge devices use unified memory, where CPUs, GPUs, and NPUs share the same DRAM.
@@
Table 1: Resource utilization (CPU, GPU, and memory) when offloading model layers to the GPU in llama.cpp. As more layers are offloaded, memory consumption increases significantly compared to CPU-only inference.
@@ 3.1 Model
Figure 4: Nanomind hardware design and PCB layout. (a) Block diagram of hardware components: an RK3566 SoC, a PMU IC for power monitoring, and four 4 GB LPDDR4x memory modules in parallel; (b) front view of PCB design; (c) back view of PCB design.
@@ 3.2 Software–Hardware Coordination
NPU Most mobile NPUs only support static computation graphs, meaning that any change in input shape requires recompiling the firmware—an impractical step for resource-constrained devices. To address this limitation, as highlighted in llm.npu (Xu et al., 2025), we offload the vision encoder to the NPU and preprocess all images by compressing and resizing them to a fixed resolution, ensuring consistent input shapes throughout inference. Rockchip’s NPU driver (RKNN) (Linux, 2025) provides native support for running models such as CLIP (Radford et al., 2021), SigLip (Zhai et al., 2023), and Whisper.cpp, achieving higher speed than open-source implementations. We deploy SigLip on the SoC’s NPU rather than the GPU primarily for performance reasons: the official RKNN driver delivers a more efficient execution environment, making the vision encoder substantially faster on the NPU. In contrast, mapping the LLM to the NPU is less practical due to its static shape restriction—prompt lengths vary dynamically, and any change in input shape requires recompiling the model. While padding inputs to a fixed maximum context length could bypass this issue, it would lead to severe inefficiency by wasting both compute cycles and memory bandwidth. In addition to the official SDK, insights from community resources—such as technical blogs and forums Devices —were instrumental in navigating RKNN conversion and optimizing operator mappings, helping us maximize NPU efficiency.
GPU Our inference framework builds on llama.cpp, retaining the ggml (GGUF) model format while extending it with a customized backend to support heterogeneous edge accelerators. Using GGUF as a unified format allows Nanomind to leverage a wide range of open-source quantized models. To further improve efficiency on resource-constrained devices, we incorporate OpenCL-based GPU kernels enhanced with linear attention and fused dequant-GEMM operations for W4A16 quantization (4-bit weights, FP16 activations). To handle sequences efficiently on resource-constrained devices, we replace standard quadratic attention with linear attention. This kernelized, streaming variant maintains running summaries of past keys and values, updating them as new tokens arrive and computing outputs via a single matrix–vector pass—avoiding the costly $T{\times}T$ score matrix. The design integrates with our W4A16 stack for fast inference. We also implement a fused dequant-GEMM OpenCL kernel that unpacks and rescales int4 weights in-register within the GEMM loop, followed immediately by FP16 FMAs. This fusion eliminates intermediate buffers and memory passes, turning each byte into useful MACs—critical for mobile GPUs without INT8 tensor cores. The kernel uses tiled vector loads, scale tables in constant/LDS memory, and an epilogue that can fuse bias and activation, with FP16/FP32 accumulators for stability. Together, these optimizations reduce memory traffic and latency while preserving accuracy.
@@
Token-Aware Buffer Management To enable efficient token flow across accelerators, Nanomind introduces the Token-Aware Buffer Manager (TABM)—a lightweight runtime module on the CPU and the core of dynamic workload offloading (Figure 3). TABM manages a shared ring buffer pool in unified DRAM and directly streams tokens between the NPU (producer) and GPU (consumer), achieving true zero-copy transfer. This design eliminates redundant memory movement, reduces CPU overhead, and cuts latency while sustaining a high-throughput token pipeline. By tracking buffer states with a simple state machine and using lightweight synchronization, TABM also smooths producer–consumer mismatches and provides scheduling signals for higher-level control.
@@ 3.3 Hardware Design
Parallel LPDDR4x Memory: To address memory bandwidth limitations in small-form-factor devices—particularly important for memory-bound LLM workloads—we integrate four LPDDR4x memory channels in parallel, significantly boosting effective memory throughput. The CPU orchestrates unified memory management across accelerators, maintaining full control over resource recycling. In particular, it governs memory clearance and buffer reuse for both GPU and NPU workloads, ensuring efficient utilization of the shared memory space.
@@ 4.1 Resource Usage
Figure 6: Throughput (tokens/s) and Latency (s) across different hardware platforms running Qwen2-VL-2B-Instruct with 4-bit quantization. “O-Pi”refers to the Orange Pi 5 Ultra, while “Nano” denotes the Nvidia Jetson Nano. Nanomind (Own) applies dynamic offloading, using FP16 for the vision encoder and W4A16 for the LLM base model.
@@ 4.2 Different Combinations of Hybrid Quantization
Figure 7 presents a comparison of different quantization strategies and module decoupling configurations. Each legend label follows the format Module–Quantization, where em-” denotes the embedding layer, vis-” refers to the vision encoder (ViT), and dec-” indicates the language decoder (Qwen2-0.5B). fp16” represents 16-bit floating-point precision, while “q4f16” refers to 4-bit weight quantization combined with fp16 activations. We evaluate these configurations across multiple benchmarks, including MMBench, MMLU, MME, and InfoVQA. The results show that when the VLM is decomposed and each module—such as the ViT and LLM—is executed independently on different accelerators, the accuracy on vision-related tasks is predominantly determined by the precision of the ViT. This highlights the importance of allocating higher bitwidth or computational resources to the vision encoder when optimizing for multimodal performance under constrained hardware.
4.3 Power consumption and Hours to use
@@ Appendix A Appendix
GPUs excel at tensor operations like matrix addition and multiplication due to their high parallelism. When a GPU is available, tensors can be designated with GGML_BACKEND_GPU, enabling ggml_compute_forward() to offload computations. The process starts by transferring key tensors (K, Q, V) to GPU memory, where mathematical operations are executed while the CPU manages the overall flow. Instead of storing results in the data pointer, they remain in GPU memory until the final operation is completed, at which point the result tensor is copied back to CPU memory.

### v3

@@ Abstract
Large Multimodal Models (LMMs) are inherently modular, consisting of vision and audio encoders, projectors, and large language models. Yet, they are almost always executed monolithically, which underutilizes the heterogeneous accelerators (NPUs, GPUs, DSPs) in modern SoCs and leads to high end-to-end latency. In this paper, we present Nanomind, a hardware–software co-design inference framework for Large Multimodal Models (LMMs) that breaks large models into modular “bricks” (vision, language, audio, etc.) and maps each to its ideal accelerator. The key insight is that large models can be broken into modular components and scheduled to run on the most appropriate compute units. It performs module-level dynamic offloading across accelerators on unified-memory SoCs. By combining customized hardware design, system-level scheduling, and optimized low-bit computation kernels, we demonstrate our framework with a compact, battery-powered device capable of running LMMs entirely on-device. This prototype functions as a self-contained intelligent assistant that requires no network connectivity, while achieving higher throughput and superior power efficiency under strict resource constraints. The design further bypasses CPU bottlenecks and reduces redundant memory usage through token-aware buffer management and module-level coordination. Our system outperforms existing implementations in resource efficiency, cutting energy consumption by 42.3% and GPU memory usage by 11.2%. This enables a battery-powered device to run LlaVA-OneVision-qwen2-05B with a camera for nearly 20.8 hours.
@@ 1 Introduction
To overcome these challenges, we introduce Nanomind, the first fully on-device inference framework that decomposes large multimodal models into modular, independently executable components and dynamically offloads each to its optimal compute unit—GPU, NPU, or CPU. Nanomind is built through a tightly integrated software–hardware co-design. We demonstrate it using a custom battery-powered prototype device (Figure 11). With this hardware platform and system-level implementation, Nanomind outperforms mainstream frameworks running on commodity off-the-shelf devices. Our contribution lies in a SW/HW co-design approach at the inference-system level, where we develop a series of system and software optimizations rather than modifying model algorithms.
We also design an event-driven On-Demand Cascade Inference Pipeline as shown in Fig. 2. Only the minimal output needed—such as a text string or an embedding vector—is retained and passed to the next stage. This results in a lightweight, domino-like chain of execution.
@@
  • Custom Hardware–Software Co-Design. On the hardware side, we use a commodity RK3566 SoC with integrated GPU and NPU, maximized memory bandwidth with in-parallel LPDDR4x modules, and add a dedicated power management unit (PMU) for real-time energy monitoring. On the software side, we implement custom 2-bit, 4-bit, and 8-bit GEMM kernels tailored to our hardware, along with an offloading scheduler and drivers to accelerate quantized tensor operations on both GPU and NPU.
@@
By employing these efforts, a tiny device can efficiently operate LLMs and LMMs (LlaVa Liu et al. (2023b; a), Qwen2-VL series Bai et al. (2023); Wang et al. (2024b)) within constrained hardware resources by directly offloading workloads to the on-device accelerators based on power and memory usage. This approach enhances inference performance and significantly reduces power consumption. Although our hardware prototype is implemented on RK3566—and we also tested key components on RK3588 (Orange Pi 5)—our framework is not tied to any specific SoC. Modern edge and mobile SoCs (e.g., Apple Silicon, Qualcomm 6590, recent RK and MediaTek chips) typically integrate multiple accelerators, often including both an NPU and a GPU, and sometimes a DSP, making cross-accelerator execution directly applicable. Our work lays the foundation for bringing LLMs to resource-constrained environments, enabling the development of responsive, power-efficient, and intelligent systems. It opens the door to democratizing LMM deployment on small devices, transforming how we interact with AI in everyday settings.
@@ 2.2 On-Device Inference Systems and Frameworks
Inefficiencies in llama.cpp While llama.cpp provides layer-wise offloading, its workload distribution is inefficient on small devices, particularly under unified memory. Although computation can be split between the CPU and GPU, GPU execution still depends on CPU-managed data transfers, increasing memory overhead during inference. Figure 10 in the Appendix illustrates this workflow, with further discussion in Section A.1. When a GPU is available, tensors can be assigned the GGML_BACKEND_GPU flag, allowing ggml_compute_forward() to offload computations to the GPU. This involves transferring key tensors from CPU memory, while the CPU must continuously write to buffers and maintain separate memory allocation, leading to additional overhead. This type of framework enables LLM deployment on edge devices but follows server-side designs with separate CPU and GPU memory. In contrast, modern edge devices use unified memory, where CPUs, GPUs, and NPUs share the same DRAM.
@@
Table 1: Resource utilization (CPU, GPU, and memory) when offloading model layers to the GPU in llama.cpp to illustrate it. Offloading more layers substantially increases memory usage relative to CPU-only inference.
@@ 3.1 Model
Figure 4: Nanomind hardware design and PCB layout. (a) Block diagram of hardware components: an RK3566 SoC, a PMU IC for power monitoring, and LPDDR4x memory modules in parallel; (b) front view of PCB design; (c) back view of PCB design.
@@ 3.2 Software–Hardware Coordination
NPU Most mobile NPUs only support static input shape, meaning that any change in input shape requires recompiling the firmware—an impractical step for resource-constrained devices. To address this limitation, we offload the vision encoder to the NPU and pre-process all images by compressing and resizing them to a fixed resolution, ensuring consistent input shapes during inference. Rockchip’s RKNN driver (Linux, 2025) and Qualcomm’s AI Hub (Qualcomm, 2025) provide native support for models such as CLIP (Radford et al., 2021) and SigLip (Zhai et al., 2023), offering higher performance than open-source implementations. We deploy SigLip on the NPU rather than the GPU for performance reasons: the official RKNN driver provides a significantly more efficient execution environment, making the vision encoder much faster on the NPU Rockchip . In contrast, running the LLM on the NPU is impractical due to its static-shape requirement—prompt lengths vary at runtime, Therefore, we offload only the Vision Encoder/ViT to the NPU. All input images are pre-processed to a fixed resolution of $448\times 736$ (Qwen2-VL) or $384\times 384$ (Llava-onevision), and multi-frame inputs are merged using a simple average temporal pooling operation. In addition to the official SDK, insights from community resources—such as technical blogs and forums Devices —were instrumental in navigating RKNN conversion and optimizing operator mappings, helping us maximize NPU efficiency.
GPU Our inference kernel on GPU builds on llama.cpp, retaining the ggml (GGUF) model format while extending it with a customized backend to support heterogeneous edge accelerators. Using GGUF as a unified format allows Nanomind to leverage a wide range of open-source quantized models. To further improve efficiency on resource-constrained devices, we incorporate OpenCL-based GPU kernels enhanced with linear attention and fused dequant-GEMM operations for W4A16 quantization (4-bit weights, FP16 activations). To handle sequences efficiently on resource-constrained devices, we replace standard quadratic attention with linear attention. This kernelized, streaming variant maintains running summaries of past keys and values, updating them as new tokens arrive and computing outputs via a single matrix–vector pass—avoiding the costly $T{\times}T$ score matrix. The design integrates with our W4A16 stack for fast inference. We also implement a fused dequant-GEMM OpenCL kernel that unpacks and rescales int4 weights in-register within the GEMM loop, followed immediately by FP16 FMAs. This fusion eliminates intermediate buffers and memory passes, turning each byte into useful MACs—critical for mobile GPUs without INT8 tensor cores. The kernel uses tiled vector loads, scale tables in constant/LDS memory, and an epilogue that can fuse bias and activation, with FP16/FP32 accumulators for stability. Together, these optimizations reduce memory traffic and latency while preserving accuracy.
@@
Token-Aware Buffer Management To enable efficient token flow across accelerators, Nanomind introduces the Token-Aware Buffer Manager (TABM)—a lightweight runtime module on the CPU and the core of dynamic workload offloading (Figure 3). TABM manages a shared ring buffer pool in unified DRAM and directly streams tokens between the NPU (producer) and GPU (consumer), achieving true zero-copy transfer. This design eliminates redundant memory movement, reduces CPU overhead, and cuts latency while sustaining a high-throughput token pipeline.
@@ 3.3 Hardware Design
Parallel LPDDR4x Memory: To address the memory-bound nature of LLM workloads on small-form-factor devices, we enhance the effective bandwidth utilization of RK3566’s LPDDR4x subsystem. Although the chip uses four DDR schedulers that multiplex into a single 32-bit controller, our coordinated CPU–GPU–NPU buffer management reduces contention and redundant transfers, improving overall memory efficiency for LLM inference.
@@ 4.1 Resource Usage
Figure 6: Throughput (tokens/s) and end-to-end latency (s) for Qwen2-VL-2B-Instruct on the InfoVQA Mathew et al. (2022) dataset across different hardware platforms. “O-Pi” refers to the Orange Pi 5 Ultra, and “Jetson” to the NVIDIA Jetson Nano. Nanomind uses cross-accelerator dynamic offloading, with FP16 for the vision encoder and W4A16 for the LLM.
@@ 4.2 Different Combinations of Hybrid Quantization
To illustrate the trade-off between quantizations and performance, Figure 13 in the Appendix compares various quantization strategies and module-decoupling configurations. Legend labels follow the format Module–Quantization, where em- denotes the embedding layer, vis- the vision encoder (ViT), and dec- the language decoder (Qwen2-0.5B). fp16 indicates 16-bit floating point, and q4f16 represents 4-bit weight quantization with fp16 activations. We evaluate these configurations on MMBench, MMLU, MME, and InfoVQA. As shown in Figure 13, when the VLM is decomposed and each module (e.g., the ViT and LLM) runs on different accelerators, accuracy on vision-related tasks is largely determined by the ViT’s precision.
4.3 Breakdown of System-Level Performance
To analyze improvements beyond end-to-end performance, we conduct a system breakdown evaluation to quantify the contribution of each component in our design.
Zero-Copy TABM Architecture vs. Traditional Copy-Based Buffering TABM’s ring-buffer and zero-copy design primarily reduce memory usage and CPU overhead by eliminating the need for CPU-managed buffer writes when transferring embeddings. We evaluated memory usage (GB) and CPU utilization (%) in Figure 7(a). Compared with llama.cpp’s memory copy and layer-wise offloading, TABM achieves lower memory usage and significantly reduced CPU load during embedding transfer.
Encoder Throughput: NPU vs. GPU vs. CPU To measure vision encoder performance across computation units, we benchmark SigLip (Zhai et al., 2023) (from LLaVA-OneVision-Qwen2 (HF, 2025)) and ArcFace (Deng et al., 2022) on the NPU, GPU, and CPU, reporting FPS under continuous image input. All images inputs to SigLip are resized to $384\times 384$, matching its training resolution to avoid performance degradation, and ArcFace uses the authors’ evaluation dataset (Guo & Deng, 2017). Figure 7(b) shows that both SigLip and ArcFace run significantly faster on the NPU, as the RKNN is optimized for vision tasks, giving the NPU a clear FPS advantage over the GPU.
(a) Zero-Copy TABM: Memory and CPU
(b) Encoder Throughput: NPU vs. GPU vs. CPU
(c) Fused dequant GEMM kernel vs Baselines
Figure 7: System Breakdown Performance. a) Zero-Copy TABM vs. Traditional Direct Copy and Offloading used on llama.cpp. b) Encoder Throughput (FPS) comparison of the SigLip ViT encoder (from LLaVA-OneVision (Liu et al., 2024a; HF, 2025)) and the ArcFace encoder (Deng et al., 2022) across RKNN NPU, Mali GPU, and CPU. c) Throughput Comparison (Tok/s): Nanomind ’s custom GEMM kernels (GPU-only LLM decoding), llama.cpp, MLC-LLM, and PowerInfer-2 on the Orange Pi 5 (RK3588) and RubikPi (Dragonwin QCS6490) while running Qwen2-1.5B-W8A8.
Fused dequant-GEMM Kernel vs. Existing Frameworks The fused dequant–GEMM kernel runs on the SoC GPU for LLM decoding. Because PowerInfer-2 and MLC-LLM currently support only LLM, we compare all frameworks in a text-only setting using the Qwen2-1.5B-W8A8 model. Figure 7(c) shows that our fused dequant–GEMM kernel achieves the highest throughput (tok/s), with PowerInfer-2 close behind. MLC-LLM on the RubikPi 3 (Qualcomm QCS6490) performs worse, likely due to weaker OpenCL support on Qualcomm GPUs rather than limitations of MLC-LLM or the hardware. This experiment evaluates only GPU-side decoding and does not involve cross-accelerator inference or buffer management.
Figure 8: Energy–Latency Trade-off Across Three Power Modes. The curve illustrates how the system adapts to the battery level ($B$). (1) In the Unconstrained State, parallel acceleration delivers low latency at higher power. (2) In the Proportional State, the system linearly throttles frame rate and memory bandwidth as $B$ decreases, producing a continuous latency–power trade-off trajectory. (3) In the Critical State, the system transitions to the low-power On-Demand Cascade pipeline.
4.4 Quantitative Power Consumption Analysis
To explore the energy–latency trade-offs across power modes during long-running workloads, Figure 8 shows how Nanomind adapts to different battery levels in a real-world smart headband deployment. As illustrated in Figure 12 in the Appendix, we built a battery-powered prototype and conducted a week-long study with five users, collecting extensive traces that were used to produce Figure 8.
@@ Appendix A Appendix
Most of the open-source frameworks—such as llama.cpp—were designed for desktops and servers with separate CPU–GPU memory, requiring repeated parameter copies from DRAM to GPU memory. Although later adapted for edge devices, they inherit assumptions from these architectures. Modern mobile SoCs use unified memory, where CPU, GPU, and NPU share the same DRAM. Applying legacy designs directly leads to inefficiencies, as accelerators must coordinate access to shared memory, making new system-level optimizations necessary.
In llama.cpp, GPUs accelerate tensor operations such as matrix multiplication through high parallelism. When a GPU backend (GGML_BACKEND_GPU) is enabled, ggml_compute_forward() offloads supported operators to the GPU. During execution, key tensors (e.g., K, Q, V) are transferred from host memory to GPU memory, where the associated computations are performed while the CPU orchestrates control flow. Intermediate results stay in GPU memory, and only the final output tensor is copied back to CPU memory once the operation completes.
@@ A.3 Measurement and Datasets
A.4 Real-World Deployment of Nanomind on a Headband
To study user experience, energy efficiency, and latency under real-world usage with variable interaction patterns, we built a headband-based demo device running Nanomind. Users wore the device and interacted with it through natural language.
(a) Head Band with Nanomind
(b) Headband with Nanomind
A.5 Different Combinations of Hybrid Quantization
The results show that when the VLM is decomposed and each module—such as the ViT and LLM—is executed independently on different accelerators, the accuracy on vision-related tasks is predominantly determined by the precision of the ViT. This highlights the importance of allocating higher bitwidth or computational resources to the vision encoder when optimizing for multimodal performance under constrained hardware.
A.6 Adaptation Across Different SoCs
Table 2 summarizes the theoretical deployment support of Nanomind across different SoCs. Our current implementation is fully realized only on custom RK3566 and RK3588 hardware with integrated PMU. Support for the Qualcomm Dragonwing QCS6490 is still under development, and our evaluations for this platform are conducted on the Rubik Pi 3.
SoC NPU GPU DSP Supported? Notes
RK3588 $\checkmark$ $\checkmark$ $\checkmark$ Directly supported
QCS6490 DSP-based NPU $\checkmark$ $\checkmark$ In-Progress Directly supported
Apple M2 $\checkmark$ $\checkmark$ Partial Closed-source
Mali-only SoC $\checkmark$ Very Limited CPU–GPU Coordinate
Table 2: SoC support.